# Optimised neutron yield calculations from ($\alpha,n$) reactions with the modified SOURCES4 code

### v1

@@ 5 Conclusions
The optimised approach for calculating neutron yields using SOURCES4 code shows a good agreement with experimental data, proving the importance of using reliable cross-section data where available. In cases where experimental data are not available for a specific isotope, nuclear reaction codes such as TALYS, EMPIRE, or databases like JENDL provide a good alternative. The choice of models for the cross-sections is confirmed by comparing calculated neutron yields with alpha beam data to ensure accuracy and reliability. This approach ensures that the most accurate and recent information is used for neutron production calculations, which is very important for applications in rare event physics experiments. By increasing the number of discrete nuclear levels considered in TALYS 1.96, the modified code now accounts for all possible excited states for the final state nucleus, as available in RIPL-3.
@@ 6 Acknowledgements
V.K. and P.K. would like to thank the Science and Technology Facilities Council (STFC, UK) and the University of Sheffield for financial support. For M.P. this work was performed with the financial support of the Romanian Program PNCDI III, Programme 5, Module 5.2 CERN-RO, under contract no. 04/2022.

### v2

@@ 5 Conclusions
The optimised approach for calculating neutron yields using SOURCES4 code shows a good agreement with experimental data, proving the importance of using reliable cross-section data where available. In cases where experimental data are not available for a specific isotope, nuclear reaction codes such as TALYS, EMPIRE, or databases like JENDL provide a good alternative. The choice of models for the cross-sections is confirmed by comparing calculated neutron yields with alpha beam data and measurements of neutron yields from radioactive decay chains to ensure accuracy and reliability. This approach ensures that the most accurate and recent information is used for neutron production calculations, which is very important for applications in rare event physics experiments. By increasing the number of discrete nuclear levels considered in TALYS 1.96, the modified code now accounts for all possible excited states for the final state nuclei, as available in RIPL-3.
The modified libraries of cross-sections and branching ratios used in this work are available at [80]. The full SOURCES4 source code cannot be shared due to licensing restrictions but the users who obtained the code from the Nuclear Energy Agency [81] are welcome to request the updated version from mihaela.parvu@unibuc.ro or v.kudryavtsev@sheffield.ac.uk.
@@ 6 Acknowledgements
V.K. and P.K. would like to thank the Science and Technology Facilities Council (STFC, UK, ST/W000547/1, ST/S003398/1, ST/Y004841/1) and the University of Sheffield for financial support. For M.P. this work was performed with the financial support of the Romanian Program PNCDI III, Programme 5, Module 5.2 CERN-RO 04/2022 and CERN-RO/CDI/2024_001.